1. Sell or Process Further

Dakota Coffee Company produces Columbian coffee in batches of 7,500 pounds. The standard quantity of materials required in the process is 7,500 pounds, which cost $6.00 per pound. Columbian coffee can be sold without further processing for $9.80 per pound. Columbian coffee can also be processed further to yield Decaf Columbian, which can be sold for $11.60 per pound. The processing into Decaf Columbian requires additional processing costs of $6,300 per batch. The additional processing will also cause a 5% loss of product due to evaporation.

a. Prepare a differential analysis dated December 11 on whether to sell regular Columbian (Alternative 1) or process further into Decaf Columbian (Alternative 2). If required, use a minus sign to indicate a loss.

c. Determine the price of Decaf Columbian that would cause neither an advantage nor a disadvantage for processing further and selling Decaf Columbian. Round your answer to two decimal places.
$ per pound
Answer #1
Decaf Columbian units 7125 =7500*(1-5%)
a
Differential Analysis
Sell Regular Columbian (Alt. 1) or Process Further into Decaf Columbian (Alt. 2)
December 11
Sell Process Differential
Regular Further into Effect
Columbian Decaf on Income
(Alternative 1) Columbian (Alternative 2)
(Alternative 2)
Revenues 73500 82650 9150
Costs -45000 -51300 -6300
Income (Loss) 28500 31350 2850
c
Price of Decaf Columbian 11.20 per pound =11.60-(2850/7125)
Workings:
Sell Process
Regular Further into
Columbian Decaf
(Alternative 1) Columbian
(Alternative 2)
Revenues =7500*9.80 =7125*11.60
Costs =7500*6 =45000+6300